OpenStudy (anonymous):

a hot bowl of soups cools according to Newtons law of cooling. its temperature (in degrees Fahrenheit) at time t is given by T(t)=68+144e^-0.4t, where t is given in minutes. A. What was the initial temperature of soup? B. What is the temperature of the soup after 15 minutes? C. How long after serving is the soup 125 degrees Fahrenheit?

OpenStudy (cwrw238):

initial temperature of the soup occurs when t = 0 so plug in t = o into the formula to get the answer to A.

OpenStudy (anonymous):

what the function is T(t)=68+144e^-0.04

OpenStudy (cwrw238):

for B plug in t =15

OpenStudy (anonymous):

i just dont know what to do with e

OpenStudy (cwrw238):

for A: T(0) = 68 + 144e^0 whats the value of e^0 ?

OpenStudy (cwrw238):

what is the value of any number to the power 0?

OpenStudy (cwrw238):

3^0 = 2^0 = 10 ^0 x^0 = ? etc all = 1

OpenStudy (cwrw238):

so 144 e^0 = 144 * 1 and t(0) = 16 + 144 = 212 thats the answer to A

OpenStudy (anonymous):

ok so for be i get